# First Light and Reionization Epoch Simulations (FLARES) - XV: The physical properties of super-massive black holes and their impact on galaxies in the early Universe

### v2

@@ 3 Blackhole Properties
In this section we explore the physical properties of SMBHs and their host galaxies in FLARES. As previously noted (in §2.2), a consequence of the simulation resolution and modelling choices (i.e. SMBH seed mass) is that we can only be confident in the properties of SMBHs with masses $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$, which we consider ”resolved”. Consequently, we focus our attention on these systems though we do explore predictions at lower-masses.
@@ 3.2 SMBH density and mass Function
Figure 3 shows the SMBH mass function predicted by FLARES, again from $z=10\to 5$. The mass function also drops by 10$\approx$ 10$\times$ 10 × from $M_{\bullet}=10^{7}\to 10^{8}\ {\rm M_{\odot}}$, largely independent of redshift. There is tentative evidence of a steeper drop at $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$, but this is complicated by the small sample size and thus large statistical uncertainty. Figure 3 also shows the SMBH mass function split by bolometric luminosity. Of note here is that the most massive SMBHs are not necessarily the most luminous; this is discussed in more detail in §3.3.2. The total mass function shown in Figure 3 is provided in Table 1 and electronically.
Figure 3: The evolution of the SMBH mass function predicted by FLARES at $5\leq z\leq 10$. The solid coloured lines in both panels show the total mass function (i.e. for all SMBHs). The shaded region around this provides an estimate of the uncertainty using the $68\%$ Poisson confidence interval. The thin coloured dashed, dot-dashed, and dotted lines on the left-hand panel denote the mass functions of SMBHs with $L_{\rm bol}/{\rm erg\ s^{-1}}<10^{45}$, $10^{45}<L_{\rm bol}/{\rm erg\ s^{-1}}<10^{46}$, and $L_{\rm bol}/{\rm erg\ s^{-1}}>10^{46}$ respectively. Shown on the right-hand panels are recent observational constraints and model predictions. The observations Matthee et al. (2023) at $z\sim 5$ and He et al. (2023) at $z\approx 4$ in the $z=5$ panel. Models predictions include ASTRID, Bluetides, EAGLE, Illustris, Simba, TNG100, and TNG300.
@@ 3.2.1 Observational Constraints
As a key physical distribution function the SMBH mass function has been the focus of considerable observational study.
@@
He et al. (2023) recently combined SDSS observations with a fainter Hyper Suprime-Cam selected sample with spectroscopic follow-up to study the SMBH mass function at $z\approx 4$. This samples ranges from $M_{\bullet}=10^{7.5-10.5}\ {\rm M_{\odot}}$ and bolometric luminosities $10^{45.5-47.5}\ {\rm erg\ s^{-1}}$. The broad line SMBH mass function at $z=4$ observed by He et al. (2023) is shown alongside our predictions in Figure 3. These constraints are consistent with our total mass function function at $\approx 10^{9}\ {\rm M_{\odot}}$ but fall short at lower-masses. However, the He et al. (2023) mass function is measured at $z=4$, not at $z=5$ where the FLARES predictions lie. Extrapolating the FLARES $z=5$ SMBH mass function to $z=4$ would suggest a density increase of around 0.5 dex. Secondly, the He et al. (2023) mass function only includes unobscured (i.e. type 1) and active SMBH and thus provides only a lower-limit on the true mass function. However, He et al. (2023) also attempt to constrain the total mass function, making corrections for obscuration and in-active SMBHs. The density of $M_{\bullet}=10^{8-9}\ {\rm M_{\odot}}$ BHs is 10100$\sim$ 10-100$\times$ 10 - 100 × larger than density of active broad line SMBHs, with the range sensitive to the modelling assumptions. The upper-end of this correction would elevate the total mass function above our predicted mass function across the full mass-range. While highly uncertain this suggests that the FLARES predictions are compatible with these observations.
JWST is now enabling a similar approach at higher-redshift and lower luminosities. Several studies (Harikane et al., 2023; Matthee et al., 2023; Maiolino et al., 2023b) have already employed this method to infer SMBH masses, and in the case of Matthee et al. (2023) the broad-line SMBH mass function. The observational constraints of Matthee et al. (2023) are shown in Figure 3 and, at first glance, provide a good match to our predictions for all SMBH. However, for the same reasons described above, the Matthee et al. (2023) mass function will provide a lower-limit on the total mass function. This suggests possible tension between our predictions and Matthee et al. (2023) and possibly Matthee et al. (2023) and He et al. (2023).
This comparison highlights the issues involved in directly comparing predictions from simulations with observations. These issues can be lessened, though not eliminated, by instead comparing bolometric luminosity functions (which we do in §3.3.2). However, this results in discarding useful information (the line-widths). An alternative approach is to forward model the SMBHs to predict observational quantities that can be analysed in the same way as observational samples. This is a current focus of work with results expected in a companion work (Wilkins et al. in-prep).
@@ 3.2.2 Model comparisons
The right-hand panel of Figure 3 also shows predictions from several other cosmological hydrodynamical simulations including Astrid, Bluetides, EAGLE, Illustris, Simba, TNG100, and TNG300. Figure 3 immediately demonstrates the power of the FLARES approach. Only the $(400/h\ {\rm Mpc})^{3}$ Bluetides simulation mass function extends as far as FLARES, and only then to $z=7$ where the simulation stopped. For all the other simulations FLARES significantly extends the mass range probed. As expected, the FLARES predictions closely match the EAGLE predictions (where they overlap) but are also similar to Illustris, Simba, TNG100, and Bluetides (at $z=7$, less so at $z>7$). The agreement with ASTRID and TNG300 is weaker with both predicting a lower normalisation than found in FLARES. A thorough comparison between many of these models is presented in Habouzit et al. (2022a) at low-redshift and Habouzit et al. (2022b) at high-redshift.
@@ 3.2.3 Environmental Dependence
One of the strengths of FLARES is its ability to probe the effect of environment on galaxy formation at high-redshift. In Figure 4 we show how the total number and mass function of SMBHs varies between simulations and thus environment, $\log_{10}(1+\delta_{14})\approx-0.3\to 0.3$.
These figures reveal, unsurprisingly, that SMBHs in FLARES are extremely biased. Virtually all of the $>10^{7}\ {\rm M_{\odot}}$ SMBHs simulated in FLARES are in the most extreme regions. While there are $\approx 700$ $>10^{7}\ {\rm M_{\odot}}$ SMBHs across the FLARES simulations at $z=5$, only $\approx 10$ are in regions with $\delta_{14}\leq 0.0$. The strong environmental dependence of SMBHs predicted by FLARES raises the issue of cosmic variance affecting observational surveys (e.g. Thomas et al., 2023).
Figure 4: Environmental dependence of the number (top panel) and mass function (bottom panel) of SMBHs predicted by FLARES at $z=5$. The top panel shows the number of robust $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs in each simulated region as a function of the region over-density $\delta_{14}$. Dotted lines indicate the behaviour if $N\propto\delta+1$. The bottom panel shows the SMBH mass function for individual simulations, coloured coded by the region over-density $\delta_{14}$. Also shown here in black is the composite SMBH mass function found by combining all forty simulations with the appropriate weighting. The horizontal grey line denotes the density corresponding to a single object in an individual FLARES simulation.
@@ 3.3 Growth
As described in §2.2 SMBHs in the EAGLE model grow through both accretion (§2.2.2) and mergers (§2.2.4). We defer an exploration of SMBH mergers in FLARES to a work in preparation (Liao et al. in-prep) and focus here on growth through accretion, including making predictions for the (accretion disc) accretion rates and bolometric luminosities of SMBHs in FLARES. As explained in §2.2.3, since we assume a fixed radiative efficiency (in our case $\epsilon_{\rm r}=0.1$) the bolometric luminosity simply scales with the accretion disc accretion rate, allowing us to explore them interchangeably.
@@ 3.3.1 Correlation with SMBH mass
We begin, in Figure 5, by showing predictions for the relationship between SMBH mass ($M_{\bullet}$) and the accretion rate, bolometric luminosity (top–panel) and Eddington ratio $\lambda$ (bottom–panel). As noted previously we can only be confident in our predictions of SMBHs with $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$. Since a large fraction of SMBHs are accreting at the Eddington limit this suggests our predictions for the bolometric luminosities of SMBHs are, conservatively, complete above the Eddington luminosity of a $10^{7}\ {\rm M_{\odot}}$ SMBH, i.e. $\approx 10^{45}\ {\rm erg/s}$ ($\approx 2.5\times 10^{11}\ {\rm L_{\odot}}$).
SMBHs in FLARES exhibit a wide range of accretion rates at fixed mass, extending up to our imposed maximum ($1/h$$\times$ the Eddington rate). For resolved SMBHs the median Eddington ratio ($\lambda$) is around $0.01$ below which it gradually drops such that there are very few galaxies with $\lambda<10^{-6}$. While the number of SMBHs with ratios higher than the median also drops, they “pile-up” at the Eddington luminosity due to the imposed limiter. The binned median Eddington ratio (denoted by the dashed line in the bottom–panel of Figure 5) drops by around $0.5$ dex $M_{\bullet}>10^{7}-10^{9}\ {\rm M_{\odot}}$. Figure 5 also shows the binned median Eddington rate but weighted by the accretion rate (solid). As would be expected this is biased towards higher accretion rates, except at the highest masses where the small numbers lead to convergence. Thus, while the typical Eddington ratios are $\sim 0.01$ most accretion (and thus most energy is produced) is SMBHs with much higher Eddington ratios.
@@ 3.3.2 Bolometric Luminosity Function
45.1 -5.4 -5.75 -6.63 -5.75 -5.16 -4.6
45.3 -6.75 -6.68 -6.17 -5.63 -4.76 -4.34
45.5 -6.9 -6.09 -6.38 -5.63 -5.83 -5.04
45.7 - - -6.75 -6.05 -5.81 -5.36
45.9 - -7.18 -7.28 -6.71 -7.04 -6.51
46.1 - - -7.45 -6.2 -6.34 -6.36
46.3 - -7.22 -5.42 -7.45 -6.29 -4.98
Next, in Figure 6, we explore predictions for the SMBH bolometric luminosity function. Due to our conservative completeness limit there are only a relatively small number of objects with luminosities above this limit, resulting in noisier predictions than the SMBH mass function. Nevertheless, this analysis reveals a clear evolution from $z=10\to 5$ with the density of SMBHs with $L_{\rm bol}>10^{45}\ {\rm erg\ s^{-1}}$ increasing by around a factor $100$.
In Figure 6, we also compare our predictions to two recent observational studies. First we compare against the “free” and “polished” variants of the quasar bolometric luminosity function presented in Shen et al. (2020) at $z=5$ and $z=6$. These luminosity functions are derived by converting multiple observations of monochromatic quasar luminosity functions in the UV, optical, hard X-ray, and mid-IR into bolometric luminosity functions and finding the best fit. In the “free” variant the fitting of $\phi_{\star}$ is left free at each redshift. However, at high-redshift there is considerable observational uncertainty. In the “polished” variant $\phi_{\star}(z)$ is assumed to be linear with observations at $z=0.4-3.0$ used to define the relationship between $z$ and $\phi_{\star}$. While both variants provide excellent agreement with the FLARES predictions at $z=5-6$ it is worth noting that $L_{\rm bol}=10^{45-46}\ {\rm erg\ s^{-1}}$ is at the limit of the Shen et al. (2020) constraints.
@@
One observational possibility for this discrepancy is potential contamination from host galaxy light or cosmic variance due to the relatively small volumes surveyed by JWST so far. Both issues will be overcome with increasing sample sizes and spectroscopic completeness. In terms of a theoretical explanation, beyond a fundamental issue with the model, there is some modelling flexibility in our predictions. First, here we use the instantaneous accretion rates, however, if we use accretion rates averaged over longer timescales we can obtain different results. This is explored in more detail in Appendix A. While the bright-end ($>10^{45.5}\ {\rm erg\ s^{-1}}$) of the LF remains immune to the choice of timescale (see Figure 13) we observe that the faint end can be enhanced by averaging over a longer timescale, bringing our predictions up to the lower luminosity constraints of Kokorev et al. (2024) and Greene et al. (2023). Second, here we assume a radiative efficiency of $\epsilon_{\rm r}=0.1$; increasing this to $\epsilon_{\rm r}=0.2$, as done by some other simulations, would shift the bolometric luminosity function by 0.3 dex to higher luminosities, again improving the agreement with Kokorev et al. (2024) and Greene et al. (2023).

### v3

@@ 3 Blackhole Properties
In this section we explore the physical properties of SMBHs and their host galaxies in FLARES. As previously noted (in §2.2), a consequence of the simulation resolution and modelling choices (i.e. SMBH seed mass) is that we can only be confident in the properties of SMBHs with masses $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$, which we consider “resolved”. Consequently, we focus our attention on these systems though we do explore predictions at lower-masses.
@@ 3.2 SMBH density and mass Function
Figure 3 shows the SMBH mass function predicted by FLARES, again from $z=10\to 5$ highlighting its evolution. Figure 3 also includes an estimate of the statistical uncertainty in each bin based on $68\%$ Poisson confidence interval. However, this is the minimum possible uncertainty based simply on the total number of SMBHs in each mass bin. In the context of FLARES this is likely a significant underestimate since mean-density simulation have small numbers of SMBHs (and thus individually large statistical uncertainties) but high weighting. The mass function also drops by an order of magnitude from $M_{\bullet}=10^{7}\to 10^{8}\ {\rm M_{\odot}}$, largely independent of redshift. There is tentative evidence of a steeper drop at $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$, but this is complicated by the small sample size and thus large statistical uncertainty. The total mass function shown in Figure 3 is provided in Table 1 and electronically.
Figure 3: The evolution of the SMBH mass function predicted by FLARES at $5\leq z\leq 10$. The solid coloured lines in both panels show the total mass function (i.e. for all SMBHs). The shaded region around this provides an estimate of the statistical uncertainty using the $68\%$ Poisson confidence interval. Also shown on each panel are models predictions from ASTRID, Bluetides, EAGLE, Illustris, Simba, TNG100, and TNG300.
@@ 3.2.1 Observational Constraints
He et al. (2023) recently combined SDSS observations with a fainter Hyper Suprime-Cam selected sample with spectroscopic follow-up to study the SMBH mass function at $z\approx 4$. This samples ranges from $M_{\bullet}=10^{7.5-10.5}\ {\rm M_{\odot}}$ and bolometric luminosities $10^{45.5-47.5}\ {\rm erg\ s^{-1}}$. These constraints are consistent with our total mass function function at $\approx 10^{9}\ {\rm M_{\odot}}$ but fall significantly below our predictions at lower-masses. However, the He et al. (2023) mass function is measured at $z=4$, not at $z=5$ where the FLARES predictions lie. Extrapolating the FLARES $z=5$ SMBH mass function to $z=4$, based on the previous evolution, would suggest a density increase of around 0.5 dex. Secondly, the He et al. (2023) mass function only includes unobscured (i.e. type 1) and active SMBH and thus provides only a lower-limit on the true mass function. However, He et al. (2023) also attempt to constrain the total mass function, making corrections for obscuration and in-active SMBHs. The density of $M_{\bullet}=10^{8-9}\ {\rm M_{\odot}}$ BHs is 10100$\sim$ 10-100$\times$ 10 - 100 × larger than density of active broad line SMBHs, with the range sensitive to the modelling assumptions. The upper-end of this correction would elevate the total mass function above our predicted mass function across the full mass-range.
JWST is now enabling a similar approach at higher-redshift and lower luminosities. Several studies (Harikane et al., 2023; Matthee et al., 2023; Maiolino et al., 2023b) have already employed this method to infer SMBH masses, and in the case of Matthee et al. (2023) the broad-line AGN SMBH mass function. The raw observational constraints of Matthee et al. (2023) provide a good match to our predictions for all SMBH. However, for the same reasons described above, the Matthee et al. (2023) mass function will provide a lower-limit on the total mass function suggesting that the true (completeness corrected) mass function will lie above our predictions. However, this also suggests possible tension between Matthee et al. (2023) and He et al. (2023).
An alternative observational test is to compare against the bolometric luminosity function since this, in principle, will suffer fewer completeness issues. However, this has its own complications. This is discussed in §3.3.3 and we defer a discussion of potential modelling changes until there.
@@ 3.2.2 Model comparisons
Figure 3 also shows predictions from several other cosmological hydrodynamical simulations including ASTRID, Bluetides, EAGLE, Illustris, Simba, TNG100, and TNG300. This Figure immediately demonstrates the power of the FLARES approach - only the very-large $(400/h\ {\rm Mpc})^{3}$ volume Bluetides simulation mass function extends as far as FLARES, and only then to $z=7$ where the simulation stopped. For all the other simulations FLARES significantly extends the mass range probed. As expected, the FLARES predictions closely match the EAGLE predictions (where they overlap) but are also similar (i.e. $<0.5$ dex difference) to Illustris, Simba, TNG100, and Bluetides (at $z=7$, less so at $z>7$). The agreement with ASTRID and TNG300 is significantly weaker with both predicting $\sim 1$ dex fewer massive SMBHs than FLARES at $z=5$ with the disagreement widening at higher-redshift.
The differences seen here reflect the different ways in which the seeding, dynamics, growth, and feedback are modelled by the different simulations. A thorough comparison between many of these models is presented in Habouzit et al. (2022a) at low-redshift and Habouzit et al. (2022b) at high-redshift. However, unfortunately, since each model [^1] Except FLARES and EAGLE, since FLARES is based on a variant of the EAGLE model. has several differences in both the SMBH and wider physics it is impossible to decisively identify the source of differences. A curiosity however is the large difference between ASTRID and FLARES/EAGLE; ASTRID seeds lower-mass halos, using a comparable mass seed [^2] ASTRID assumes a probabilistic approach for the seed masses, drawing from a power-law distribution $M_{\bullet,{\rm seed}}/h^{-1}\ {\rm M_{\odot}}=[3\times 10^{4},3\times 10^{5}]$ with power-law index $n=-1$., and permits more rapid accretion but results in 10$\sim$ 10$\times$ 10 × fewer massive BHs at $z=5$. This suggests that the AGN feedback, or some other feature of the model, is key here. A priority for the next phase of FLARES is to systemically explore the impact of the SMBH modelling choices, including the choice of parameters. This should provide a much clearer understanding of the key factors driving the growth of SMBHs in the early Universe.
@@ 3.2.3 Environmental Dependence
One of the strengths of FLARES is its ability to probe the effect of environment on galaxy formation at high-redshift. In Figure 4 we show how the total number (top-panel) and mass function (bottom panel) of SMBHs varies between simulations and thus environment, $\log_{10}(1+\delta_{14})\approx-0.3\to 0.3$.
These figures reveal, unsurprisingly, that SMBHs in FLARES are significantly biased. Virtually all of the resolved ($>10^{7}\ {\rm M_{\odot}}$) SMBHs simulated in FLARES are found in the most extreme regions. While there are $\approx 700$ $>10^{7}\ {\rm M_{\odot}}$ SMBHs across the FLARES simulations at $z=5$, only $\approx 10$ are in regions with $\delta_{14}\leq 0.0$. This bias is even more extreme for more massive SMBHs, with SMBHs of $>10^{8}\ {\rm M_{\odot}}$ only lying in the most extreme over-densities that we simulate. The SMBH mass function of our most over-dense regions ($\log_{10}(1+\delta_{14}\approx 0.25$) is around $1$ (1.5) ${\rm dex}$ higher at $M_{\bullet}=10^{7}\ {\rm M_{\odot}}$ ( $M_{\bullet}=10^{8}\ {\rm M_{\odot}}$) higher than the weighted average.
This then suggests that massive SMBHs (and bright AGN) should pinpoint over-dense regions. This has recently been explored by Eilers et al. (2024) who used JWST/NIRCam observations to explore the environments of 4 luminous ($\sim 10^{46}\ {\rm erg\ s^{-1}})$ quasars at $z\geq 6$. Eilers et al. (2024) found that these quasars do indeed on average trace large over-densities but they do not necessarily trace the rarest and highest density peaks. However, it is not immediately straightforward to compare the Eilers et al. (2024) results with our predictions. This is because the Eilers et al. (2024) quasars are brighter than any AGN in FLARES at $z\geq 6$, they use a smaller search radius, and use [Oiii] emitting galaxies.
Figure 4: Environmental dependence of the number (top panel) and mass function (bottom panel) of SMBHs predicted by FLARES at $z=5$. The top panel shows the number of robust $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs in each simulated region as a function of the region over-density $\delta_{14}$. Dotted lines indicate the behaviour if $N\propto\delta+1$. The bottom panel shows the SMBH mass function for every individual simulation with at least one SMBH in the mass range
, colour coded by the region over-density $\delta_{14}$. Also shown here in black is the composite SMBH mass function found by combining all forty simulations with the appropriate weighting. The horizontal grey line denotes the density corresponding to a single object in an individual FLARES simulation.
@@ 3.3 Growth
As described in §2.2 SMBHs in the EAGLE model grow through both accretion (§2.2.2) and mergers (§2.2.4), albeit with mergers producing only a small fraction of the overall growth of resolved SMBHs in FLARES. We defer an exploration of SMBH mergers in FLARES to a work in preparation (Liao et al. in-prep) and focus here on growth through accretion, including making predictions for the accretion rates and bolometric luminosities of SMBHs in FLARES.
3.3.1 Accretion rate definitions
By default the accretion rate associated with each SMBH in a particular snapshot is that calculated in the most recent time-step. However, these instantaneous accretion rates are numerically noisy. Fortunately, within FLARES SMBH accretion rates are recorded at every time-step, permitting an exploration of accretion rate histories.
As an alternative to the instantaneous accretion rate we can average the accretion over a longer timescale ( $t_{\rm avg}/{\rm Myr}=10-200$) in an attempt to reduce the numerical noise. Figure 5 shows the ratio of the instantaneous accretion rate to the accretion rate averaged over the preceding 10 Myr for SMBHs at $z=5$. This reveals that the majority of galaxies have average accretion rates larger than their instantaneous rate. However, the total accretion integrated over all galaxies is similar as would be expected for a large number of SMBHs.
Figure 5: The ratio of the instantaneous accretion rate to the accretion rate averaged over the preceding 10 Myr for SMBHs at $z=5$. The solid line denoted the median in $M_{\bullet}$ bins. Individual BHs are colour coded by their Eddington ratio ($\lambda$) based on the instantaneous accretion rate.
While the total accretion integrated across all SMBHs is similar irrespective of the timescale there is an impact on the shape of the accretion rate distribution function, or bolometric luminosity function. Figure 6 shows the bolometric luminosity function calculated assuming the different accretion timescales $\in\{10,20,50,100,200\}\ {\rm Myr}$. While the density of very-luminous ($>10^{46}\ {\rm erg\ s^{-1}}$) SMBHs remains relatively unchanged, the number of lower luminosity, but still bright ($<10^{45}\ {\rm erg\ s^{-1}}$) SMBHs increases. This is due to many SMBHs with very low instantaneous accretion rates having significantly higher averaged rates as demonstrated by Figure 5.
Since the instantaneous accretion rates are subject to numerical noise we decide to adopt the 10 Myr averaged accretion rate as our fiducial measure in the remainder of our analysis. It is important to stress however that the luminosities of real AGN are well established to vary on timescales much shorter than this, or possible to resolve in any cosmological simulation. This is a clear limitation of this type of analysis.
@@ 3.3.2 Correlation with SMBH mass
We next explore, in Figure 7, predictions for the relationship between SMBH mass ($M_{\bullet}$) and the accretion rate, bolometric luminosity (top–panel) and Eddington ratio $\lambda_{\rm Edd}$ (bottom–panel). As noted previously we can only be confident in our predictions of SMBHs with $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$. Since a large fraction of SMBHs are accreting at the Eddington limit this suggests our predictions for the bolometric luminosities of SMBHs are, conservatively, complete above the Eddington luminosity of a $10^{7}\ {\rm M_{\odot}}$ SMBH, i.e. $\approx 10^{45}\ {\rm erg/s}$ ($\approx 2.5\times 10^{11}\ {\rm L_{\odot}}$).
SMBHs in FLARES exhibit a wide range of accretion rates at fixed mass, extending close to the imposed maximum ($1/h$$\times$ the Eddington rate). Few SMBHs have accretion rates at the maximum since (as discussed in §3.3.1) we average over the last 10 Myr instead of using instantaneous accretion rates. For resolved SMBHs the median Eddington ratio is around $0.1$ below which it gradually drops such that there are very few galaxies with $\lambda<10^{-4}$. While the number of SMBHs with ratios higher than the median also drops, they “pile-up” at towards the imposed limit. The binned median Eddington ratio (denoted by the dashed line in the bottom–panel of Figure 7) drops by around $1.0$ dex $M_{\bullet}>10^{7}-10^{9}\ {\rm M_{\odot}}$. Figure 7 also shows the binned median Eddington rate but weighted by the accretion rate (solid). As would be expected this is biased towards higher accretion rates, except at the highest masses where the small numbers lead to convergence. Thus, while the typical Eddington ratios are $\sim 0.1$ most accretion (and thus most energy is produced) is SMBHs with much higher Eddington ratios ($\sim 1$).
@@ 3.3.3 Bolometric Luminosity Function
44.5 -6.53 -5.1 -4.67 -4.68 -4.2 -4.01
44.75 -6.44 -5.07 -5.29 -4.88 -4.55 -4.19
45.0 -5.5 -6.32 -6.04 -5.32 -4.73 -4.44
45.25 -6.22 - -6.44 -5.27 -5.37 -4.98
45.5 - - -6.62 -6.32 -4.71 -4.98
45.75 - - -6.72 -6.39 -6.04 -5.53
46.0 - - - -6.25 -6.41 -5.83
46.25 - - - - - -5.07
As noted in §2.2.3, since we assume a fixed radiative efficiency (in our case $\epsilon_{\rm r}=0.1$) the bolometric luminosity simply scales with the accretion disc accretion rate, allowing us to explore them interchangeably. Since the bolometric luminosity is observationally accessible in Figure 8, we explore predictions for the SMBH bolometric luminosity function. Due to our conservative completeness limit there are only a relatively small number of objects with luminosities above this limit, resulting in noisier predictions than the SMBH mass function. Nevertheless, this analysis reveals a clear evolution from $z=10\to 5$ with the density of SMBHs with $L_{\rm bol}>10^{45}\ {\rm erg\ s^{-1}}$ increasing by around a factor $100$.
In Figure 8, we also compare our predictions to two recent observational studies. First we compare against the “free” and “polished” variants of the quasar bolometric luminosity function presented in Shen et al. (2020) at $z=5$ and $z=6$. These luminosity functions are derived by converting multiple observations of monochromatic quasar luminosity functions in the UV, optical, hard X-ray, and mid-IR into bolometric luminosity functions and finding the best fit. In the “free” variant the fitting of $\phi_{\star}$ is left free at each redshift. However, at high-redshift there is considerable observational uncertainty. In the “polished” variant $\phi_{\star}(z)$ is assumed to be linear with observations at $z=0.4-3.0$ used to define the relationship between $z$ and $\phi_{\star}$. At $z=5$ the normalisation at the bright-end ($L\sim 10^{45.5-46}\ {\rm erg\ s^{-1}}$) is in good agreement with the Shen et al. (2020) model; however, at fainter luminosities FLARES predicts more SMBHs. At $z=6$ FLARES predicts a similar shape and normalisation to the “polished” variant across the full luminosity range. However, it is important to note that the FLARES predictions cover the low-luminosity limit of the Shen et al. (2020) model.
@@
[…]
There are also several modelling factors which may explain this tension and also possible the potential tension in the mass function.
First, as noted in §3.3.1, we average accretion rates over a 10 Myr timescale to reduce numerical noise. However, the emission from SMBHs is well established to often vary on much shorter timescales. While this would impact the shape of the bolometric luminosity it seems extremely unlikely that it could account for the significant tension at the highest redshifts.
It’s possible that these tensions could also be explained by changes to the model. Properly evaluating the impact of changes to the model is beyond the scope of this work, since any change will manifest in complex ways, impacting the SMBH mass and luminosity function, as well as the properties of host galaxies. For example, a boost to the accretion at one time may lead to a subsequent suppression due to feedback. This type of exploration is a key priority for the next phase of the FLARES project.
However, it is nevertheless useful to consider the impact of some simple changes. First, we could assume a larger radiative efficiency $\epsilon_{\rm r}$. At fixed accretion rate this would boost the bolometric luminosity; however, it would, assuming the same Eddington limiter, also reduce SMBH growth through the impact on Eddington rate (Equation 2) and SMBH mass growth (Equation 5). Since the Bondi accretion rate is proportional to $M_{\bullet}^{2}$ such a change would likely suppress the luminosities of the most luminous SMBHs further exasperating the tension. Alternatively, we could relax the Eddington limiter. Since the most luminous SMBHs are growing at (or close to) the limit this would likely boost the growth and luminosities of the most luminous SMBHs. While this may create better agreement at the highest redshifts, it would likely create a new tension at lower redshift ($z\approx 5$) where the agreement is currently good. Another option is to increase the seed mass; this would likely manifest in larger masses and accretion later one; again removing tension at the highest redshifts, albeit at the expense of adding tension at lower-redshift.